OpenStudy (anonymous):

If y = 2 when x = –4, find x when y = 20.

OpenStudy (anonymous):

a.y=-1/2x;-40 b. y=1/2x;40 c. y=-x;-20 d. y=-1/2x;-79/2

OpenStudy (aravindg):

Well how does y vary with x ? directly or inversely?